OpenStudy (anonymous):

what is equation of the line passing through -3 -9 and parallel to line y=x+6 in standard form x+y=-6 x+y=6 x-y=-6 x-y=6

OpenStudy (anonymous):

x-y=6

OpenStudy (anonymous):

what is the vertex of the graph of y=|3x-12|+5 (-4,-5) (4,-5) (-4,5) (4,5)

OpenStudy (anonymous):

y-5=3x-12 and y-5=12-3x solve these two u will get (4,5)